OpenStudy (anonymous):

find the largest of three consecutive integers such that 3 times the sum of the first and third integer is equal to 30 more than 4 times the second integer

OpenStudy (ddcamp):

So we have 3 integers, a,b, and c. Using these integer variables, can you write equations that describe the problem?

OpenStudy (anonymous):

3(a+c)=30+4(b)?

OpenStudy (ddcamp):

Yup!

OpenStudy (thomas5267):

Make use of the fact that the three integers are consecutive.

OpenStudy (ddcamp):

Since they're consecutive, try writing all of the variables in terms of the largest one (the one you're solving for).

OpenStudy (anonymous):

umm.. would it be 3(a+ c=3)= 30+4(b+2)????

OpenStudy (anonymous):

no i don't think that's it

OpenStudy (ddcamp):

Let's try redefining a and b: If a, b, and c are consecutive, then that means: a = c - 2 b = c - 1 That make sense?

OpenStudy (anonymous):

yes that does

OpenStudy (ddcamp):

Try replacing a and b in your original equation with (c-1) and (c-2).

OpenStudy (anonymous):

3(c-1+c)=30+4(c-1)

OpenStudy (anonymous):

i ment 3(c-2+c)=30+4(c-1)

OpenStudy (ddcamp):

Good! Now, just solve for c.

OpenStudy (anonymous):

c=16

OpenStudy (ddcamp):

Yup! Good job.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

thank you so much